# GAN Based Near-Field Channel Estimation for Extremely Large-Scale MIMO Systems

### v1

@@ I Introduction
To tackle this issue, some near-field channel estimation schemes have been proposed for extremely large-scale antenna array (ELAA) based systems. The authors of [16] proposed a subarray-wise channel estimation approach to estimate the near-field non-stationary channel with low complexity. However, the 2-dimentional distance-angle plane in [16] was uniformly divided into multiple grids to construct the near-field transform matrix, which fails to reduce the correlation among the array steering vectors. To exploit the sparsity of the near-field channel, a polar-domain representation based channel estimation method was proposed in [17] by simultaneously considering both the angle and distance information. However, the power of nonzero elements for the near-field channel may leak to other elements when the polar-domain sparsity is considered, which leads to a power leakage effect [18]. The authors of [19] proposed a hybrid-field channel estimation method to estimate the near-field and far-field channel parameters by using the OMP algorithm. The authors of [20] proposed a model based deep learning (DL) method for XL-MIMO systems by incorporating the sparsity dictionary into one neural network, but the performance needs to be improved.
@@
The vast majority of existing near-field channel approaches for XL-MIMO systems usually introduces the channel sparse representation scheme and utilize the CS based method such as the OMP algorithm to estimate the near-field channel. The sparse representation approach leads to high computational complexity because the 2-dimentional distance-angle plane for the near-field channel needs to be partitioned into grids [20]. The sparse representation scheme also assumes that the signal falls precisely on these grids, thus limiting the channel estimation accuracy [21]. Furthermore, the near-field XL-MIMO channel model adopted in the aforementioned works [16, 17, 18, 19, 20, 21] can not describe the line-of-sight (LoS) path component of the near-field XL-MIMO channel accurately. Hence, these existing near-field channel estimation approaches are not capable of accurately estimating the XL-MIMO near-field channel, especially when both the transmitter and receiver are equipped with ELAAs [22, 23]. To address this problem, the authors of [22] proposed a near-field XL-MIMO channel model by simultaneously considering LoS and non-line-of-sight (NLoS) path components. A two stage channel estimation approach was also proposed in [22] by separately estimating the LoS and NLoS path components. However, separately estimating the LoS and NLoS path components may lead to high computational complexity. Moreover, the NLoS path components are estimated in [22] by using an OMP based algorithm, which limits the channel estimation performance.
@@
The authors of [24] have proposed generative adversarial networks (GANs) as a framework in which the generator and the discriminator are adversarially trained to learn an adversarial loss. Recently, GANs have achieved remarkable success in image generation such as sketch generation [25] and image-to-image translation [26]. The key success of GANs is to learn an adversarial loss that allows the model to generate images similar to the real images. Since the channel matrices can be treated as two-channel images, the problem of the channel estimation in wireless communications can be regarded as an image-to-image translation problem. Hence, the GAN and its variants can be exploited to solve the channel estimation problem. Recently, the GAN and its variants such as GANs [27, 28], conditional GAN (CGAN) [29, 30] and Wasserstein GAN (WGAN)[31] have been successfully used to solve the problem of far-field channel estimation. For instance, the authors of [31] integrated an improved WGAN with a CGAN to enhance the channel estimation performance of the traditional least-square (LS) estimation scheme. However, GAN architectures in the aforementioned works are not designed for near-field channel estimation problems in XL-MIMO communication systems.
@@
Notations: We use bold-face lower-case letter $\mathbf{a}$ and upper-case letter $\mathbf{A}$ to denote a vector and a matrix, respectively. $\left\|\mathbf{A}\right\|_{2}$ represents the $l_{2}$-norm of matrix $\mathbf{A}$. $\left(\cdot\right)^{H}$, $\left(\cdot\right)^{-1}$, $\mathbb{E}\left[\cdot\right]$, and $\log\left(\cdot\right)$ denote conjugate transpose, inverse, expectation, and logarithmic operators, respectively. $\mathcal{C}\mathcal{N}\left(\mu,\sigma^{2}\right)$ denotes circularly symmetric complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. $\mathcal{U}\left(-a,a\right)$ represents the uniform distribution on $\left(-a,a\right)$. $\mathbf{I}$ denotes an identity matrix.
@@ II System Model
In this section, we first introduce the signal model of near-field XL-MIMO systems, and then describe the adopted near-field XL-MIMO channel model.
@@ II-B The Adopted Near-Field XL-MIMO Channel Model
Let $d$, $\alpha_{l}$, and $L$ denote the antenna spacing, the complex gain at the $l$-th path, and the number of NLoS path components, respectively. Let $\theta_{\mathrm{r},l}$ and $\theta_{\mathrm{t},l}$ denote the angles of the $l$-th path at the receiver and the transmitter, respectively. $d_{l}^{\mathrm{t}}$ ($d_{l}^{\mathrm{r}}$) denotes the distance between the center of the antenna array of the transmitter (receiver) and the $l$-th scatterer. Let $\mathbf{b}\left(\theta_{\mathrm{r},l},d_{l}^{\mathrm{r}}\right)$ and $\mathbf{b}\left(\theta_{\mathrm{t},l},d_{l}^{\mathrm{t}}\right)$ be the near-field array response vectors at the receiver and the transmitter under the spherical wave assumption, respectively. Let $d_{l}^{\mathrm{t}}\left(u\right),u=1,2,\cdots,N_{\mathrm{t}},$ be the distance between the $u$-th element on the transmitter antenna array and the $l$-th scatterer as $d_{l}^{\mathrm{t}}\left(u\right)=\sqrt{\left(d_{l}^{\mathrm{t}}\right)^{2}+d^{2}\delta_{u}^{2}-2dd_{l}^{\mathrm{t}}\delta_{u}\sin\theta_{\mathrm{t},l}}$, where $\delta_{u}=\small{\frac{2u-N_{\mathrm{t}}-1}{2}}$. Let $d_{l}^{\mathrm{r}}\left(v\right),v=1,2,\cdots,N_{\mathrm{r}},$ denote the distance between the $v$-th element on the receiver antenna array and the $l$-th scatterer as $d_{l}^{\mathrm{r}}\left(v\right)=\sqrt{\left(d_{l}^{\mathrm{r}}\right)^{2}+d^{2}\delta_{v}^{2}-2dd_{l}^{\mathrm{r}}\delta_{v}\sin\theta_{\mathrm{r},l}}$, where $\delta_{v}=\small{\frac{2v-N_{\mathrm{r}}-1}{2}}$. Then, the XL-MIMO near-field channel [22] only containing NLoS path components can be expressed as
@@
$\frac{1}{r_{v,u}}$ $=\sqrt{\left(r\cos\theta-d_{2}\sin\varphi\right)^{2}+\left(r\sin\theta+d_{2}\cos\varphi-d_{1}\right)^{2}}$ (7)
@@
According to [22] and [17], the advanced RD (ARD) and RD are defined as $\small{\frac{4D_{\mathrm{t}}D_{\mathrm{r}}}{\lambda}}$ and $\small{\frac{2\left(D_{\mathrm{t}}+D_{\mathrm{r}}\right)^{2}}{\lambda}}$, where $D_{\mathrm{t}}=\small{\frac{\lambda N_{\mathrm{t}}}{2}}$ and $D_{\mathrm{r}}=\small{\frac{\lambda N_{\mathrm{r}}}{2}}$. $D_{\mathrm{t}}$ and $D_{\mathrm{r}}$ denote the apertures of the antenna array at the transmitter and the receiver, respectively. The boundary between the existing XL-MIMO near-field channel model and the mixed LoS/NLoS XL-MIMO near-field channel model adopted in this paper is provided in Fig. 2. When the distance $r$ is smaller than the ARD, the receiver is in the mixed LoS/NLoS near-field region of the transmitter, while the receiver is in the NLoS near-field region with the distance $r$ increasing to a certain extent. We also show the boundary between the existing XL-MIMO near-field channel model and the far-field MIMO channel model in Fig. 2. The receiver is in the far-field region when the distance $r$ is larger than the AD. In this paper, we mainly focus on the channel estimation in the mixed LoS/NLoS near-field region.
@@ III-A Initial Channel Estimation
where $\mathbf{G}_{\mathrm{L}}=\mathbf{W}^{H}\left(\mathbf{WW}^{H}\right)^{-1}$ and $\mathbf{G}_{\mathrm{R}}=\mathbf{Q}^{H}\left(\mathbf{QQ}^{H}\right)^{-1}$. After the initial channel estimation, the initially estimated channels $\mathbf{H}_{\mathrm{IE}}\in\mathbb{C}^{N_{\mathrm{r}}\times N_{\mathrm{t}}}$ can be obtained. Then, the initially estimated channels instead of the received signals are fed into the proposed IE-Pix2pix as the conditional input. Finally, the proposed IE-Pix2pix outputs the estimated near-field XL-MIMO channels through the mapping relationship
@@ III-B Objective Functions
The channel estimation problem is treated as an image-to-image translation problem because the channel matrices can be treated as two-channel images. Therefore, we are able to exploit the GAN and its variants to address the problem of channel estimation. To estimate the channel matrices, GANs are used to learn the mapping from the random noise $\mathbf{Z}$ to the true channel matrices. The noise $\mathbf{Z}$ is usually sampled from a prior distribution $p_{Z}$ (e.g., uniform distribution and Gaussian distribution). However, this mapping learned by GANs may be unstable and the learning capacity of GANs is usually not good [30, 31] for channel estimation tasks. To tackle this issue, both the random noise $\mathbf{Z}$ and the conditional input are input into the CGAN to generate channel matrices with a specific property. As one of the GAN variants, Pix2pix is utilized to learn the mapping from the conditional input to the true channel matrices, which is similar to the CGAN. However, no random noise $\mathbf{Z}$ is directly fed into the Pix2pix. Although Pix2pix may produce deterministic outputs, it can still learn the mapping from the conditional input to the channel matrices [32]. In this paper, the proposed IE-Pix2pix based on the Pix2pix is developed to estimate the XL-MIMO near-field channels more accurately by using the initially estimated channel matrices as the conditional input. In the following, we take the GAN based near-field channel estimation as an example to describe the objective functions and structures of the GAN and its variants.
@@
respectively. In the proposed IE-Pix2pix, $G\left(\mathbf{H}_{\mathrm{IE}}\right)$ denotes the mimic channel samples produced by the generator, while $D\left(G\left(\mathbf{H}_{\mathrm{IE}}\right),\mathbf{H}_{\mathrm{IE}}\right)$ or $D\left(\mathbf{H},\mathbf{H}_{\mathrm{IE}}\right)$ represents the output of the discriminator. During the training, the generator learns to minimize each of the elements in $\log\left(1-D\left(G\left(\mathbf{H}_{\mathrm{IE}}\right),\mathbf{H}_{\mathrm{IE}}\right)\right)$ as seen in (16), which denotes that the generator aims to maximize each of the elements in $D\left(G\left(\mathbf{H}_{\mathrm{IE}}\right),\mathbf{H}_{\mathrm{IE}}\right)$. That is to say, the generator learns to make each of the elements in $D\left(\mathbf{H},\mathbf{H}_{\mathrm{IE}}\right)$ approach 0. On the contrary, the objective of the discriminator is to make each of the elements in $D\left(\mathbf{H},\mathbf{H}_{\mathrm{IE}}\right)$ approach 1. Therefore, the generator and the discriminator are adversarially trained to learn an adversarial loss. The objective function of the proposed IE-Pix2pix can be expressed as
@@
Finally, our objective function can be expressed as
@@
IE-Pix2pix Generator 2  $\left(\mathrm{ConvT}\right)$ 2 $3\times 3$ ReLU $1\times 1$
3  $\left(\mathrm{Conv}\right)$ 2 $3\times 3$ LReLU $1\times 1$
$4\sim 8$  $\left(\mathrm{Conv}\right)$ 64 $3\times 3$ LReLU $2\times 2$
$9\sim 12$  $\left(\mathrm{ConvT}\right)$ 64 $3\times 3$ ReLU $2\times 2$
$13$  $\left(\mathrm{ConvT}\right)$ 2 $3\times 3$ Tanh $2\times 2$
Discriminator $14\sim 16$  $\left(\mathrm{Conv}\right)$ 64 $3\times 3$ LReLU $2\times 2$
$17$   $\left(\mathrm{Conv}\right)$ 64 $3\times 3$ LReLU $2\times 1$
$18$   $\left(\mathrm{Conv}\right)$ 1 $3\times 3$ Sigmoid $1\times 1$
@@ III-C Network Architecture
The proposed IE-Pix2pix contains the generator and the discriminator. Both the generator and the discriminator are composed of convolutional layers. For the proposed IE-Pix2pix, we adopt the U-Net architecture and the patch architecture [30, 32] in the generator and the discriminator, respectively. The generator in the IE-Pix2pix mainly includes five encoder blocks and four decoder blocks. Each of the encoder blocks contains a convolutional layer (called Conv), a batch normalization layer, and a leaky-rectified linear unit (LReLU) activation layer. Every decoder block consists of a transposed convolutional layer (called ConvT), a batch normalization layer, and a rectified linear unit (ReLU) activation layer. In addition, one transposed convolutional layer and one convolutional layer are used as the first two layers of the generator, which can rescale the input of the generator into the same size as $\mathbf{H}$. When the conditional input and the near-field channel matrix are of different sizes, the conditional input can be rescaled into the same size as the near-field channel matrix by changing the strides in the first two layers of the generator. Note that a transposed convolutional layer and a convolutional layer can be treated as a upsampling layer and a downsampling layer, respectively. Filters with the size of $3\times 3$ are used for all convolutional layers. Different from the common generator, the generator in the IE-Pix2pix is an encoder-decoder CNN with skip connections. The pixel level details can be preserved by exploiting the skip connections to combine the feature maps of the decoder blocks and the encoder blocks. The tangent activation function is used as the last layer of the generator to scale the output values of the generator into the $\left[-1,1\right]$ range, which is able to avoid the mode collapse. The detailed architecture of the IE-Pix2pix is shown in Table I.

### v2

@@ I Introduction
To tackle this issue, some near-field channel estimation schemes have been proposed for extremely large-scale antenna array (ELAA) based systems. The authors of [16] proposed a subarray-wise channel estimation approach to estimate the near-field non-stationary channel with low complexity. However, the 2-dimentional distance-angle plane in [16] was uniformly divided into multiple grids to construct the near-field transform matrix, which fails to reduce the correlation among the array steering vectors. To exploit the sparsity of the near-field channel, a polar-domain representation based channel estimation method was proposed in [17] by simultaneously considering both the angle and distance information. However, the power of nonzero elements for the near-field channel may leak to other elements when the polar-domain sparsity is considered, which leads to a power leakage effect [18]. The authors of [19] proposed a model based deep learning (DL) method for XL-MIMO systems by incorporating the sparsity dictionary into one neural network, but the performance needs to be improved.
@@
The vast majority of existing near-field channel approaches for XL-MIMO systems usually introduces the channel sparse representation scheme and utilize the CS based method such as the OMP algorithm to estimate the near-field channel. The sparse representation approach leads to high computational complexity because the 2-dimentional distance-angle plane for the near-field channel needs to be partitioned into grids [19]. The sparse representation scheme also assumes that the signal falls precisely on these grids, thus limiting the channel estimation accuracy [20]. Furthermore, the near-field XL-MIMO channel model adopted in the aforementioned works [16, 17, 18, 19, 20] can not describe the line-of-sight (LoS) path component of the near-field XL-MIMO channel accurately. Hence, these existing near-field channel estimation approaches are not capable of accurately estimating the XL-MIMO near-field channel, especially when both the transmitter and receiver are equipped with ELAAs [21, 22]. To address this problem, the authors of [22] developed a hybrid-field channel model including both the near-field and far-field path components, and proposed a estimation method to respectively estimate the near-field and far-field path components. In [23], the far-field and near-field path components of the hybrid-field XL-MIMO channel were estimated by using support detection and the OMP algorithm, respectively. To accurately describe the near-field XL-MIMO channel, the authors of [24] proposed a near-field XL-MIMO channel model by simultaneously considering LoS and non-line-of-sight (NLoS) path components. A two stage channel estimation approach was also proposed in [24] by separately estimating the LoS and NLoS path components. However, separately estimating the LoS and NLoS path components may lead to high computational complexity. Moreover, the NLoS path components were estimated in [24] by using an OMP based algorithm, which limits the channel estimation performance.
@@
The authors of [25] have proposed generative adversarial networks (GANs) as a framework in which the generator and the discriminator are adversarially trained to learn an adversarial loss. Recently, GANs have achieved remarkable success in image generation such as sketch generation [26] and image-to-image translation [27]. The key success of GANs is to learn an adversarial loss that allows the model to generate images similar to the real images. Since the channel matrices can be treated as two-channel images, the problem of the channel estimation in wireless communications can be regarded as an image-to-image translation problem. Hence, the GAN and its variants can be exploited to solve the channel estimation problem. The loss functions in traditional DL based approaches are not well designed, which limits the channel estimation performance. To address this issue, the GAN and its variants such as GANs [28, 29], conditional GAN (CGAN) [30, 31], and Wasserstein GAN (WGAN)[32] have been successfully used to solve the problem of far-field channel estimation. For instance, the authors of [29] accurately estimated the channel with few pilots by optimizing the random input of the generator in the GAN according to the received signals. In [30], the received signals were fed into the CGAN as the conditional input to estimate the cascaded channels in intelligent reflecting surface assisted MIMO systems. The authors of [32] integrated an improved WGAN with a CGAN to enhance the channel estimation performance of the traditional least-square (LS) estimation scheme. The adversarial loss introduced by the GAN architecture can be used to solve the problem of loss function design in the conventional DL based schemes, which compensates for the information loss during training neural networks [31]. However, GAN architectures in the aforementioned works are not designed for near-field channel estimation problems in XL-MIMO communication systems.
@@
[…]
@@ II System Model
In this section, we first introduce the signal model of the considered near-field XL-MIMO system, and then describe the adopted near-field XL-MIMO channel model.
@@ II-B The Adopted Near-Field XL-MIMO Channel Model
Let $d$, $\alpha_{l}$, and $L$ denote the antenna spacing, the complex gain at the $l$-th path, and the number of NLoS path components, respectively. Let $\theta_{\mathrm{r},l}$ and $\theta_{\mathrm{t},l}$ denote the angles of the $l$-th path at the receiver and the transmitter, respectively. $d_{l}^{\mathrm{t}}$ ($d_{l}^{\mathrm{r}}$) denotes the distance between the center of the antenna array of the transmitter (receiver) and the $l$-th scatterer. For uniform linear arrays (ULAs), $\mathbf{b}\left(\theta_{\mathrm{r},l},d_{l}^{\mathrm{r}}\right)$ and $\mathbf{b}\left(\theta_{\mathrm{t},l},d_{l}^{\mathrm{t}}\right)$ denote the near-field array response vectors at the receiver and the transmitter under the spherical wave assumption, respectively. Let $d_{l}^{\mathrm{t}}\left(u\right),u=1,2,\cdots,N_{\mathrm{t}},$ be the distance between the $u$-th element on the transmitter antenna array and the $l$-th scatterer as $d_{l}^{\mathrm{t}}\left(u\right)=\sqrt{\left(d_{l}^{\mathrm{t}}\right)^{2}+d^{2}\delta_{u}^{2}-2dd_{l}^{\mathrm{t}}\delta_{u}\sin\theta_{\mathrm{t},l}}$, where $\delta_{u}=\small{\frac{2u-N_{\mathrm{t}}-1}{2}}$. Let $d_{l}^{\mathrm{r}}\left(v\right),v=1,2,\cdots,N_{\mathrm{r}},$ denote the distance between the $v$-th element on the receiver antenna array and the $l$-th scatterer as $d_{l}^{\mathrm{r}}\left(v\right)=\sqrt{\left(d_{l}^{\mathrm{r}}\right)^{2}+d^{2}\delta_{v}^{2}-2dd_{l}^{\mathrm{r}}\delta_{v}\sin\theta_{\mathrm{r},l}}$, where $\delta_{v}=\small{\frac{2v-N_{\mathrm{r}}-1}{2}}$. Then, the XL-MIMO near-field channel [24] only containing NLoS path components can be expressed as
@@
$r_{v,u}$ $=\sqrt{\left(r\cos\theta-d_{2}\sin\varphi\right)^{2}+\left(r\sin\theta+d_{2}\cos\varphi-d_{1}\right)^{2}}$ (7)
@@
According to [24] and [17], the advanced RD (ARD) and RD are defined as $\small{\frac{4D_{\mathrm{t}}D_{\mathrm{r}}}{\lambda}}$ and $\small{\frac{2\left(D_{\mathrm{t}}+D_{\mathrm{r}}\right)^{2}}{\lambda}}$, where $D_{\mathrm{t}}=\small{\frac{\lambda N_{\mathrm{t}}}{2}}$ and $D_{\mathrm{r}}=\small{\frac{\lambda N_{\mathrm{r}}}{2}}$. $D_{\mathrm{t}}$ and $D_{\mathrm{r}}$ denote the apertures of the antenna array at the transmitter and the receiver, respectively. The boundary between the existing XL-MIMO near-field channel model and the mixed LoS/NLoS XL-MIMO near-field channel model adopted in this paper is provided in Fig. 2. When the distance $r$ is smaller than the ARD, the receiver is in the mixed LoS/NLoS near-field region of the transmitter, while the receiver is in the NLoS near-field region with the distance $r$ increasing to a certain extent. We also show the boundary between the existing XL-MIMO near-field channel model and the far-field MIMO channel model in Fig. 2. The receiver is in the far-field region when the distance $r$ is bigger than the RD. In this paper, we mainly focus on the channel estimation in the mixed LoS/NLoS near-field region.
@@ III-A Initial Channel Estimation
where $\mathbf{G}_{\mathrm{L}}=\mathbf{W}^{H}\left(\mathbf{WW}^{H}\right)^{-1}$ and $\mathbf{G}_{\mathrm{R}}=\mathbf{Q}^{H}\left(\mathbf{QQ}^{H}\right)^{-1}$. After the IE, the initially estimated channels $\mathbf{H}_{\mathrm{IE}}\in\mathbb{C}^{N_{\mathrm{r}}\times N_{\mathrm{t}}}$ can be obtained. Then, the initially estimated channels instead of the received signals are fed into the proposed IE-Pix2pix as the conditional input. Finally, the proposed IE-Pix2pix outputs the estimated near-field XL-MIMO channels through the mapping relationship
@@ III-B Objective Functions
The channel estimation problem is treated as an image-to-image translation problem because the channel matrices can be treated as two-channel images. Therefore, we are able to exploit the GAN and its variants to address the problem of channel estimation. To estimate the channel matrices, GANs are used to learn the mapping from the random noise $\mathbf{Z}$ to the true channel matrices. The noise $\mathbf{Z}$ is usually sampled from a prior distribution $p_{Z}$ (e.g., uniform distribution and Gaussian distribution). However, this mapping learned by GANs may be unstable and the learning capacity of GANs is usually not good [31, 32] for channel estimation tasks. To tackle this issue, both the random noise $\mathbf{Z}$ and the conditional input are input into the CGAN to generate channel matrices with a specific property. As one of the GAN variants, Pix2pix is utilized to learn the mapping from the conditional input to the true channel matrices, which is similar to the CGAN. However, no random noise $\mathbf{Z}$ is directly fed into the Pix2pix. Although the Pix2pix without the noise $\mathbf{Z}$ as the input may produce deterministic outputs, it can still learn the mapping from the conditional input to the channel matrices [33]. Note that as the input of the generator in the GAN and its variants, the random noise $\mathbf{Z}$ is not the noise in the environment but the random noise sampled from a prior distribution, e.g., uniform distribution. In this paper, the proposed IE-Pix2pix based on the Pix2pix is developed to estimate the XL-MIMO near-field channels more accurately by using the initially estimated channel matrices as the conditional input. In the following, we take the GAN based near-field channel estimation as an example to describe the objective functions and structures of the GAN and its variants.
@@
respectively. In the proposed IE-Pix2pix, $G\left(\mathbf{H}_{\mathrm{IE}}\right)$ denotes the mimic channel samples produced by the generator, while $D\left(G\left(\mathbf{H}_{\mathrm{IE}}\right),\mathbf{H}_{\mathrm{IE}}\right)$ or $D\left(\mathbf{H},\mathbf{H}_{\mathrm{IE}}\right)$ represents the output of the discriminator. During the training, the generator learns to minimize each of the elements in $\log\left(1-D\left(G\left(\mathbf{H}_{\mathrm{IE}}\right),\mathbf{H}_{\mathrm{IE}}\right)\right)$ as seen in (16), which denotes that the generator aims to maximize each of the elements in $D\left(G\left(\mathbf{H}_{\mathrm{IE}}\right),\mathbf{H}_{\mathrm{IE}}\right)$. That is to say, the generator learns to make each of the elements in $D\left(\mathbf{H},\mathbf{H}_{\mathrm{IE}}\right)$ approach 0. On the contrary, the objective of the discriminator is to make each of the elements in $D\left(\mathbf{H},\mathbf{H}_{\mathrm{IE}}\right)$ approach 1. Therefore, the generator and the discriminator are adversarially trained to learn an adversarial loss. The adversarial loss for the proposed IE-Pix2pix can be expressed as
@@
To ensure the correct direction of training the generator, we combine the adversarial loss with the conventional loss function. Finally, the training process of the IE-Pix2pix can be expressed as
@@
IE-Pix2pix Generator 2  $\left(\mathrm{ConvT}\right)$ 2 $3\times 3$ ReLU $\left(1,1\right)$
3  $\left(\mathrm{Conv}\right)$ 2 $3\times 3$ LReLU $\left(1,1\right)$
$4\sim 8$  $\left(\mathrm{Conv}\right)$ 64 $3\times 3$ LReLU $\left(2,2\right)$
$9\sim 12$  $\left(\mathrm{ConvT}\right)$ 64 $3\times 3$ ReLU $\left(2,2\right)$
$13$  $\left(\mathrm{ConvT}\right)$ 2 $3\times 3$ Tanh $\left(2,2\right)$
Discriminator $14\sim 16$  $\left(\mathrm{Conv}\right)$ 64 $3\times 3$ LReLU $\left(2,2\right)$
$17$   $\left(\mathrm{Conv}\right)$ 64 $3\times 3$ LReLU $\left(2,1\right)$
$18$   $\left(\mathrm{Conv}\right)$ 1 $3\times 3$ Sigmoid $\left(1,1\right)$
@@ III-C Network Architecture
The proposed IE-Pix2pix contains the generator and the discriminator. Both the generator and the discriminator are composed of convolutional layers. For the proposed IE-Pix2pix, we adopt the U-Net architecture and the patch architecture [31, 33] in the generator and the discriminator, respectively. The generator in the IE-Pix2pix mainly includes five encoder blocks and four decoder blocks. Each of the encoder blocks contains a convolutional layer (called Conv), a batch normalization layer, and a leaky-rectified linear unit (LReLU) activation layer. Every decoder block consists of a transposed convolutional layer (called ConvT), a batch normalization layer, and a rectified linear unit (ReLU) activation layer. In addition, one transposed convolutional layer and one convolutional layer are used as the first two layers of the generator, which can rescale the input of the generator into the same size as $\mathbf{H}$. When the conditional input and the near-field channel matrix are of different sizes, the conditional input can be rescaled into the same size as the near-field channel matrix by changing the strides in the first two layers of the generator. Note that a transposed convolutional layer and a convolutional layer can be treated as a upsampling layer and a downsampling layer, respectively. Filters with the size of $3\times 3$ are used for all convolutional layers. Different from the common generator, the generator in the IE-Pix2pix is an encoder-decoder CNN with skip connections. The pixel level details can be preserved by exploiting the skip connections to combine the feature maps of the decoder blocks and the encoder blocks. The tangent activation function is used as the last layer of the generator to scale the output values of the generator into the $\left[-1,1\right]$ range, which is able to avoid the mode collapse. The detailed architecture of the IE-Pix2pix is shown in Table I. Let stride = $\left(s^{\mathrm{h}},s^{\mathrm{v}}\right)$ in Tensorflow, where $s^{\mathrm{v}}$ and $s^{\mathrm{h}}$ are the vertical and horizontal strides, respectively. Let $F_{s-1,w}$ and $F_{s-1,h}$ denote the width and the height of the input feature map in the $s$-th layer, respectively. The width and the height of the output feature map in the $s$-th transposed convolutional layer can be expressed as $F_{s,w}=\lceil F_{s-1,w}\times s^{\mathrm{h}}\rceil$ and $F_{s,h}=\lceil F_{s-1,h}\times s^{\mathrm{v}}\rceil$, respectively. The width and the height of the output feature map in the $s$-th convolutional layer can be expressed as $F_{s,w}=\lceil\frac{F_{s-1,w}}{s^{\mathrm{h}}}\rceil$ and $F_{s,h}=\lceil\frac{F_{s-1,h}}{s^{\mathrm{v}}}\rceil$, respectively.